A standard solution of hydrated Iron (II) sulphate, FeSO4.xH2O, as prepared by dissolving 5.8g of the salt in a volumetric flask making it to 250ml (250 cm3), 25ml (25 cm3) of this solution was acidified and found to react with 22.8ml (22.8 cm3) of 0.018MPotassium permanganate solution.

Calculate:

i) the molar mass of the salt.

ii) The value of x, the number of molecules of the water of crystallization.


1
Expert's answer
2021-01-28T05:04:27-0500

10FeSO4 + 2KMnO4 + 8H2SO4 → 5Fe2(SO4)3 + 2MnSO4 + K2SO4 + 8H2O

22.8 mL of 0.018 M Potassium permanganate solution:

Moles: 22.8/1000 x 0.018 = 0.00041 mol KMnO4

Moles of FeSO4: 10/2 x 0.00041 = 0.00205 mol

This amount is found in 25 mL soln

In 250 mL: 10 x 0.00205 = 0.0205 mol

i) the molar mass of the salt:

M = m/n = 5.8 g / 0.0205 mol = 282.9 g/mol

ii) The value of x, the number of molecules of the water of crystallization:

M (FeSO4) = 151.9 g/mol

282.9 – 151.9 = 131.0 refers to x H2O   

x = 131.0/18 ~ 7 molecules, where 18 – molar mass of water.